# Interface instability of two-phase flow in a three-dimensional porous medium

### v1

@@ IV.5 Beadpack characterization
Several of our experiments show anomalous front stability, including Exps. C, E, and G, which are stable experiments at low flow rates, and Exp. $Q$, which was unstable even though it was the highest flow rate studied and the most negative $\frac{d(p_{b}-p_{t})}{dt}$ (Table 2). Here we investigate whether these anomalies originate from differences in these experiments’ porous structures. We consider three main metrics: (1) recognition of small crystalline regions (crystallites), (2) local porosity distributions, and (3) radial correlation functions.
@@ V Discussion
Two independent methods were adopted for front-stability classification. Time-resolved front widths were classified as stable or unstable with a slope-over-threshold method (Sec. IV.1), indicating that fronts usually remain stable for flow rates larger than $Q_{c}\approx 8$ ml/min. Further analysis of the data using the pressure measurements and experimentally-determined front velocities, see Sec. IV.3, predicted a critical flow rate $Q_{c}=(5.5\pm 2.9)$ ml/min. Both of these independent $Q_{c}$ estimates produce equivalent values within the experimental uncertainty. Partly, this uncertainty originates from the slow variation of $d(p_{b}-p_{t})/dt$ near $Q_{c}$, see Fig. 10.
@@ V.4 Comparing with existing theory
In previous studies of drainage in quasi-2D spatially homogeneous porous media, a dimensionless number $F$ – termed the fluctuation number – has been proposed to compare the effective pressure drop developed in the fluids at the pore scale to the width of the fluctuations in the capillary-pressure thresholds to invade pore throats [25, 26, 30, 41]. Positive $F$ values indicate that the invasion front during drainage should reach a stable width, while negative values are associated with unstable fronts, in which the growth of fingers is unbounded. Besides serving as a front-stability criterion, the fluctuation number can be used to predict the front width, with $W\propto F^{-\nu/(1+\nu)}$, where $\nu$ is the critical exponent for the correlation length in 2D percolation, corresponding to $4/3$.
@@ VI Conclusion
We conducted 3D imaging experiments to assess the stability of fluid invasion fronts in a porous medium for different imposed flow rates. Above a critical flow rate $Q_{c}$, fronts stabilized at a well-defined width, whereas below this rate they destabilized. Two distinct mechanisms of front growth were visible, which we termed “pinning” and “fingering”. While fingering describes the primary instability mechanism in earlier 2D experiments, pinning describes a backward-trailing front growth which may predominate in unstable 3D invasion flows, considering the higher connectivity of 3D porous media. We estimated $Q_{c}$ using two independent methods, one based the 3D images, and another based on pressure measurements. Most of our experiments align with the $Q_{c}$ predictions, while the experiments which deviate may have been affected by the limited system size and the presence of crystalline regions in the beadpacks. Our analyses reveal several interesting consequences of trapped defending fluid behind stable invasion fronts which to our knowledge have not been emphasized in earlier studies. First, the front velocity relates non-linearly to the flow rate, and second, the relative permeability for the invading phase is not a constant, but rather has a weak dependence on the flow rate. Incorporating these dependences was found necessary to describe the observed front-velocity and pressure-measurement data within a Darcy-based framework. While our results provide useful insights into the mechanisms of front instability in 3D fluid-invasion processes, future studies will be required to further investigate invasion-front morphology and to explore the application of theories developed for 2D to the description of 3D flows.
@@ Acknowledgments
Because increasing flow rates reduce trapping [30], $S^{\prime}(Q)$ must be negative, implying that $C_{2}>0$.

### v2

@@ IV.5 Beadpack characterization
Several of our experiments show anomalous front stability, including Exps. C, E, and G, which are stable experiments at low flow rates, and Exp. $Q$, which was unstable even though it was the highest flow rate studied and the most negative $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t$ (Table 2). Here we investigate whether these anomalies originate from differences in these experiments’ porous structures. We consider three main metrics: (1) recognition of small crystalline regions (crystallites), (2) local porosity distributions, and (3) radial correlation functions.
@@ V Discussion
Two independent methods were adopted for front-stability classification. Time-resolved front widths were classified as stable or unstable with a slope-over-threshold method (Sec. IV.1), indicating that fronts usually remain stable for flow rates larger than $Q_{c}\approx 8$ ml/min. Further analysis of the data using the pressure measurements and experimentally-determined front velocities, see Sec. IV.3, predicted a critical flow rate $Q_{c}=(5.5\pm 2.9)$ ml/min. Both of these independent $Q_{c}$ estimates produce equivalent values within the experimental uncertainty. Partly, this uncertainty originates from the slow variation of $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t$ near $Q_{c}$, see Fig. 10.
@@ V.4 Comparing with existing theory
In previous studies of drainage in quasi-2D spatially homogeneous porous media, a dimensionless number $F$ – termed the fluctuation number – has been proposed to compare the effective pressure drop developed in the fluids at the pore scale to the width of the fluctuations in the capillary-pressure thresholds to invade pore throats [25, 46, 26, 30, 48]. Positive $F$ values indicate that the invasion front during drainage should reach a stable width, while negative values are associated with unstable fronts, in which the growth of fingers is unbounded. Besides serving as a front-stability criterion, the fluctuation number can be used to predict the front width, with $W\propto F^{-\nu/(1+\nu)}$, where $\nu$ is the critical exponent for the correlation length in percolation, corresponding to $4/3$ in 2D. In the absence of spatial gradients in capillary pressure thresholds, the $F$ number is effectively proportional to the difference between the Bond and Capillary numbers used in Eq. (8) as a criterion for front stability [26].
@@ VI Conclusion
We conducted 3D imaging experiments to assess the stability of fluid invasion fronts in a porous medium for different imposed flow rates. Above a critical flow rate $Q_{c}$, fronts stabilized at a well-defined width, whereas below this rate they destabilized. Two distinct mechanisms of front growth were visible, which we termed “pinning” and “fingering”. While fingering describes the primary instability mechanism in earlier 2D experiments, pinning describes a backward-trailing front growth that may predominate in unstable 3D invasion flows, considering the higher connectivity of 3D porous media. We estimated $Q_{c}$ using two independent methods, one based on the 3D images, and another based on pressure measurements. Most of our experiments align with the $Q_{c}$ predictions, while the experiments which deviate may have been affected by the limited system size and the presence of crystalline regions in the beadpacks. Our analyses reveal several interesting consequences of trapped defending fluid behind stable invasion fronts which to our knowledge have not been emphasized in earlier studies. First, the front velocity relates non-linearly to the flow rate, and second, the relative permeability for the invading phase is not a constant, but rather has a weak dependence on the flow rate. Incorporating these dependences was found necessary to describe the observed front-velocity and pressure-measurement data within a Darcy-based framework. While our results provide useful insights into the mechanisms of front instability in 3D fluid-invasion processes, future studies will be required to further investigate invasion-front morphology and to explore the application of theories developed for 2D to the description of 3D flows.
@@ Acknowledgments
Because increasing the flow rate reduces trapping [30], $S^{\prime}(Q)$ must be negative, implying that $C_{2}>0$.